Solve the following systems of equations using SUBSTITUTION to eliminate one variable. If the system has no solution, indicate that it is inconsistent; if it has infinite solutions, indicate that it is DEPENDANT; if the system has one solution, CHECK your answer.
y+5x=17
4x-3y=6

PLEASE AND THANK YOU. I AM SUPER BEHIND IN SCHOOL.

Answer by Alan3354(69443)

y+5x=17 --> y = -5x + 17
4x-3y=6
------------
Sub for y in the 2nd equation
4x-3y=6
4x - 3(-5x + 17) = 6
4x + 15x - 51 = 6
19x = 57
x = 3
-------
y = -5x + 17
y = -15 + 17
y = 2
=========